AS7000(1L) AS7000(1L) NAME as7000−crossassemblerformicrocomputers SYNOPSIS as7000[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input DESCRIPTION Theas7000commandassemblestheinputfileintoatextoutputfilerepresentingtheprogram memoryforamicrocomputer. Options -pcpuIgnoredbythisprogram. -llistfile Outputafileformatedwiththememoryaddress,data,andsourceinputlines. -ohexfile Outputthememorydataimageinaformacceptedbymostpromprogrammers. -hhexfile Sameas−o. −dSavetheintermediatedatafile(seeFILE)andaborttheexecutionattheendofprocessing. -ssymbolfile Printthesymboltablevaluesandnames,oneperlineinthespecifiedfile. FILES /usr/tmp/frtXXXXXX SEEALSO FrankensteinCrossAssemblersUsersManual(filebase.doc)Appendixforas7000Frankenstein Assembler(fileas7000.doc) NOTES Thereisonlyoneinputfile. 1